OpenStudy (koilds):

At 9:00 on Saturday morning, two bicyclists heading in opposite direction pass each other on a bicycle path. The bicyclist heading north is riding 4km/hour faster than the bicyclist heading south. At 10:15, they are 40 km apart. Find the two bicyclists' rates. A: Northbound bicyclist = 18 km/h; southbound bicyclist = 14 km/h B: northbound bicyclist = 19 km/h; southbound bicyclist = 15 km/h C: northbound bicyclist = 18 km/h; southbound bicyclist = 13 km/h D: northbound bicyclist = 16 km/h; southbound bicyclist = 10 km/h

OpenStudy (anonymous):

I found this online it might help: let v = the rate of the bicyclist heading south v + 4 = the rate of the bicyclist heading north (4 km/hour faster) d = 40 km the sum of the distances that each of the bicyclists passes t = 10:15 - 9:00 = 1 hr 15 min = 1.25 hr since the rate = distance/time follows distance = rate*time and we have v*t + (v + 4)*t = d v*1.25 + (v + 4)*1.25 = 40 by solving we find v = 14 km/hour click here to see the step by step solution of the equation http://doyourmath.com/redirect/174/3f55d74973c0ac77c742ad7033610d42?q=c%3Dsolve_algequationsolve%26v27%3Dv%2A1.25%2B%252B%2B%28v%2B%252B%2B4%29%2A1.25%2B%253D%2B40%26v28%3Dv v + 4 = 14 + 4 = 18 km/hour. The rate of the bicyclist heading south is 14 km/hour. The rate of the bicyclist heading north is 18 km/hour.

OpenStudy (koilds):

Thanks :)

OpenStudy (anonymous):

Sure ;)

TheSmartOne (thesmartone):

\(\bf\huge~~~~\color{#ff0000}{W}\color{#ff2000}{e}\color{#ff4000}{l}\color{#ff5f00}{c}\color{#ff7f00}{o}\color{#ffaa00}{m}\color{#ffd400}{e}~\color{#bfff00}{t}\color{#80ff00}{o}~\color{#00ff00}{O}\color{#00ff40}{p}\color{#00ff80}{e}\color{#00ffbf}{n}\color{#00ffff}{S}\color{#00aaff}{t}\color{#0055ff}{u}\color{#0000ff}{d}\color{#2300ff}{y}\color{#4600ff}{!}\color{#6800ff}{!}\color{#8b00ff}{!}\\\bf ~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~Made~by~TheSmartOne\) Hey there!!! Since you are new here, read this legendary tutorial for new OpenStudiers!! http://openstudy.com/study#/updates/543de42fe4b0b3c6e146b5e8 @koilds @~ You might not have noticed but you are in the "OpenStudy Feedback" subject. Hover your mouse where it says OpenStudy Feedback and find the correct subject by clicking the 'Find More Subjects' button or simply clicking the correct subject in the list that you have already studied! Alternately, you can get to the correct subject by going to this link: http://openstudy.com/study#/groups/Mathematics